ATK Tähtitieteessä 11.04.2001

*–* Ohjelma vs. komentorivi

1. IDL-proseduurit Pääohjelma

*–* .run ohjelma(.pro) *–* end

Aliohjelmat PRO aliohjelma,arg1,arg2,...,argN,keyword1=keyword1,keyword2=keyword2,... *–* Keywords: muotoa tulosta=tulosta,/tulosta,tulosta=2

Funktiot

FUNCTION funktion_nimi,arg1,arg2,...,keyword1=keyword1,keyword2=keyword2,...

2. Komentorakenteet For-silmukka

*–* Syntax: FOR i=0,10,2 DO print,i (BEGIN ... ENDFOR) If-rakenne

*–* Syntax: IF ... THEN BEGIN ... ENDIF ELSE BEGIN ... ENDELSE

While-rakenne

3. PS-tiedoston luonti

4. 2D-plotit

5. Harjoitustöiden jako